# Tight analyses of first-order methods with error feedback

### v1

@@ Abstract
Communication between agents often constitutes a major computational bottleneck in distributed learning. One of the most common mitigation strategies is to compress the information exchanged, thereby reducing communication overhead. To counteract the degradation in convergence associated with compressed communication, error feedback schemes—most notably $\mathrm{EF}$ and $\mathrm{EF^{21}}$—were introduced. In this work, we provide a tight analysis of both of these methods. Specifically, we find the Lyapunov function that yields the best possible convergence rate for each method—with matching lower bounds. This principled approach yields sharp performance guarantees and enables a rigorous, apples-to-apples comparison between $\mathrm{EF}$, $\mathrm{EF^{21}}$, and compressed gradient descent. Our analysis is carried out in a simplified yet representative setting, which allows for clean theoretical insights and fair comparison of the underlying mechanisms.
@@ 1 Introduction
The compression operator $\mathcal{C}$ is a deterministic function such that, for some $\epsilon\in[0,1)$,
for $\text{for all}\quad x\in\mathbb{R}^{d},\quad\|x-\mathcal{C}(x)\|^{2}\leqslant\epsilon\|x\|^{2}.$ (2)
@@
More recently, a variant of the classic error feedback mechanism, known as $\mathrm{EF^{21}}$ was introduced by Richtarik et al. [14], and is presented in Algorithm 3. Unlike classic error feedback, $\mathrm{EF^{21}}$ focuses on communicating a gradient estimate that is more robust to the variance observed in gradients received from different agents around the minimum of finite sum objectives. This method has since been extended in several directions [e.g. 23, 36, 37, 38].
@@
To make this comparison sharp and transparent, we adopt a deliberately simple yet representative setup: we consider smooth and strongly convex functions in the single-agent setting ($n=1$). While simple, this regime is widely recognized as a crucial stepping stone—not only for building intuition, but also as its own theoretical contribution [e.g., 32, 10]. In this context, tightness means that we identify the best possible Lyapunov function within a given class and compute the exact worst-case convergence rate over the class of problems considered.
@@
We make the following contributions:
@@
    From a methodological perspective, this paper is the first to apply the performance estimation framework to algorithms from the federated learning literature that incorporate compression schemes. By leveraging this methodology—both analytically and numerically—we pave the way for a more precise and reliable understanding of federated and distributed learning methods.
@@
    We provide a tight analysis of $\mathrm{EF}$ and $\mathrm{EF^{21}}$, and compare them with compressed gradient descent in the single-agent setting, on $L$-smooth, $\mu$-strongly convex functions. In particular, we give an analytical formula of the best possible contraction rate, by analyzing an optimal Lyapunov function within a class of candidate Lyapunov functions defined in Definition 1. Furthermore, we provide the optimal tuning for the step size in both those algorithms.
@@
    We demonstrate that those rates are achieved, proving that our analysis is tight.
@@ 2.1 Theoretical results on $\mathrm{CGD}$, $\mathrm{EF}$, $\mathrm{EF^{21}}$
In the single agent case, we leverage the equivalence between compressed gradient descent ($\mathrm{CGD}$), under Assumption 1 and the inexact gradient method with relatively bounded gradients. $\mathrm{CGD}$ corresponds to the particular case of Algorithm 1 with $n=1$, and relatively bounded gradients means that for any $x$, the oracle queried at point $x$ outputs a value $g$ such that $\|g-\nabla f(x)\|^{2}\leq\epsilon\|x\|^{2}$. Various notions of gradient approximation have been studied [48, 49, 50], and a tight analysis for relatively bounded gradients was given in [51]. Specifically, authors have shown that the inexact gradient method then enjoys tight convergence guarantees for any step size $\eta>0$, with respect to the functional residual, Euclidean norm distance to the solution and gradient norm. However, $\mathrm{CGD}$ is known to diverge when applied using stochastic gradient oracles, and to non-smooth functions [10]. Interestingly, it is also known to diverge in the multi-worker setting [24]. Studying $\mathrm{CGD}$ is important in its own right, because when the compression operator is chosen as the sign function, and the algorithm is applied in the stochastic setting (i.e., signSGD), there is a connection to Adam both in the convex [52], and non-convex setting [53].
Richtarik et al. [14] study the multi-worker setting, with (potentially) randomized compression operators. They establish a $\mathcal{O}(k^{-1})$ convergence rate on Lipschitz smooth functions, and a linear rate under the additional assumption that the functions satisfy the Łojasiewicz inequality. These results are obtained using a Lyapunov function; however, without tightness guarantees—neither for the choice of Lyapunov function, nor for the convergence rate itself. Extensions of $\mathrm{EF^{21}}$ have been proposed, including adaptations to stochastic gradients [23], and the introduction of a momentum term to improve sample complexity in the stochastic setting [36].
@@ 2.2 Performance estimation
Another relevant line of work we leverage to discover the analytical form of the Lyapunov functions lies in the field of symbolic regression, which aims to solve supervised learning tasks over the space of simple analytic expressions. Recent advances have focused on using genetic programming to search this space. Software packages have been developed in both Python and Julia, making these techniques more accessible [60].
@@ 2.3 Definitions & Notation
We have already introduced the notion of a contractive compression operator in Assumption 1. To position our contribution within the broader literature we now specify that our analysis is restricted to the setting of smooth, strongly convex functions:
@@
The function $f$ is $L$-smooth, i.e., for all $x,y\in\mathbb{R}^{d}$, we have $f(y)\leqslant f(x)+\langle\nabla f(x),y-x\rangle+\frac{L}{2}\|y-x\|^{2}.$
@@
The function $f$ is $\mu$-strongly convex, i.e., for all $x,y\in\mathbb{R}^{d}$, we have $f(y)\geqslant f(x)+\langle\nabla f(x),y-x\rangle+\frac{\mu}{2}\|y-x\|^{2}.$
@@
We formally denote :×$\mathcal{M}:\mathbb{R}^{\ell\times d}\times\mathbb{R}^{d}\times\mathcal{F}\to\mathbb{R}^{\ell\times d}\times\mathbb{R}^{d}$ a first-order method acting on a set of functions $\mathcal{F}$, for an integer $\ell\in\mathbb{N}$. Such a method, given a function $f\in\mathcal{F}$, is applied to an initial state $\xi_{0}\in\mathbb{R}^{\ell\times d}$ and iterate $x_{0}\in\mathbb{R}^{d}$, and generates a sequence $\{\xi_{k}\}_{k\geqslant 0}$ of states, and a sequence $\{x_{k}\}_{k\geqslant 0}$ of iterations. The states represent information summarizing the current point in the optimization trajectory that the algorithms may depend on beyond the current iterate—for example, error-related quantities in error feedback algorithms. The integer $\ell$ is thus typically small, from 0 to 3 in general. The specific states used in this paper are all specified in Subsection 3.1.
Definition 1 (Candidate Lyapunov functions).
@@
    (Zero at fixed-point) $\mathcal{V}(\xi,x;f)=0$ if and only if $x=x_{\star}$ and $\xi=\xi_{\star}$ and for a unique $\xi_{\star}\in\mathbb{R}^{\ell\times d}$.
@@
for some constant $\rho<1$ and for all $k\geqslant 0$, uniformly over the class $\mathcal{F}$. Finding the optimal Lyapunov function within a parameterized class, for a method $\mathcal{M}$, then amounts to solving the following problem:
@@
Note that guaranteeing contraction between step 1 and step 0, uniformly over $(\xi_{0},x_{0})$, is equivalent to guaranteeing uniform contraction between any two consecutive steps, as in (5).
@@ 3.1 Numerical performance of all methods
where all variables are defined as in Algorithm 1, Algorithm 2 and Algorithm 3, respectively.
@@
We are now ready to present the numerical results on the performance of each method. Figure 1 shows contour plots of the contraction factor for each method. That is, for a fine grid of both the step size $\eta$ in Algorithms 1, 2 and 3, and the parameter $\epsilon$ in Assumption 1, we numerically compute the value of the best possible worst-case contraction rate in terms of our class of candidate Lyapunov functions, over $\mathcal{F}_{\mu,L}$, as given by (6). A darker blue point indicates a stronger contraction $\rho^{\star}(\mathcal{M})$ (i.e., a better rate). A red point indicates that the method is non-convergent for that choice of $(\epsilon,\eta)$. We observe that, although the results are purely numerical at this stage, they indicate that $\mathrm{EF}$ and $\mathrm{EF^{21}}$ exhibit identical performance in our setting. This numerical equivalence is supported by Table 2: the maximum absolute difference between contraction factors for $\mathrm{EF}$ and $\mathrm{EF^{21}}$ is on the order of $10^{-5}$ to $10^{-7}$. This first fact is a very surprising observation. Indeed while $\mathrm{EF}$ and $\mathrm{EF^{21}}$ are known to be identical [14, see Section 4.2] in the very specific case of using a deterministic positively homogeneous and additive compression operator, $\mathrm{EF}$ and $\mathrm{EF^{21}}$ remain grounded in fundamentally different motivations at first sight: on the one hand, $\mathrm{EF}$ accounts for the errors introduced by the compression step, while on the other hand, $\mathrm{EF^{21}}$ subtracts a control variate from the gradient prior to compression. Proving that the best possible convergence rate they can obtain, for any tuning $\epsilon,\eta$, is similar (but achieved for a different Lyapunov function) was, to the best of our knowledge, never established in the literature. It thus constitutes a significant step towards better understanding their connections.
$\kappa=10.0$ $\kappa=4.0$ $\kappa=2.0$
Absolute error 1.21e-06 6.02e-07 4.34e-07
@@
$\kappa=2.0$ $\kappa=4.0$ $\kappa=10.0$
@@
Table 2: Maximum absolute difference of contraction factor for $\mathrm{CGD}$ when allowing any combination of terms in our Lyapunov function, compared to the contraction achieved by the functional residual. Same grid as Table 2.
@@
While given for a single $(\mu,L)$ in Figure 1 and Table 2, similar results hold for all values of $(\mu,L)$ that were tried numerically, and several examples are given in Appendix C, along with details of the numerical experiments, including how the regions of non-convergence were computed for these methods.
Furthermore, we tune each algorithm by picking the optimal step size for each method. We compute the rate $\inf_{\eta}\rho^{\star}(\mathcal{M}_{\eta})$, for $\mathcal{M}\in\{\mathrm{CGD},\mathrm{EF},\mathrm{EF^{21}}\}$, where $\mathcal{M}_{\eta}$ corresponds to the method with step size $\eta$. Results are shown in Figure 2 for three values of $\kappa$, namely 10, 4, and 2. In each setup, for every level of compression, $\mathrm{CGD}$ achieves a rate which is strictly better than $\mathrm{EF}$ and $\mathrm{EF^{21}}$. These results challenge the prevailing intuition that error feedback ensures convergence comparable to that of uncompressed methods, and even demonstrate that in the single agent and deterministic regime, error feedback is actually always detrimental to convergence.
Figure 2: Single row of line plots showing the performance of $\mathrm{CGD}$ (blue) and $\mathrm{EF}$ (red) as a function of the compression factor $\epsilon$, for $L=1$, and varying $\mu$. Note that each algorithm has been tuned to use an optimal step size setting.
Finally, we note that the functional residual constitutes an optimal Lyapunov function for $\mathrm{CGD}$, as shown in Table 2. This result is not particularly surprising, given that a tight analysis of the functional residual with the optimal step size was previously established for inexact gradient descent by De Klerk et al. [51]. That work also demonstrated the tightness of this rate [^1] Following the same line of reasoning as in our remark on the tightness of our Lyapunov functions in Section 5, we can then show that the functional residual is an optimal Lyapunov function for $\mathrm{CGD}$..
@@
3.2 Exact convergence rate and optimal tuning for Classic error feedback $\mathrm{EF}$
We begin by stating the main result of this section, which is a tight rate of convergence for the classic error feedback algorithm within our class of candidate Lyapunov functions.

### v2

@@ Abstract
Communication between agents often constitutes a major computational bottleneck in distributed learning. One of the most common mitigation strategies is to compress the information exchanged, thereby reducing communication overhead. To counteract the degradation in convergence associated with compressed communication, error feedback schemes—most notably $\mathrm{EF}$ and $\mathrm{EF^{21}}$—were introduced. In this work, we provide a tight analysis of both of these methods. Specifically, we find the Lyapunov function that yields the best possible convergence rate for each method—with matching lower bounds. This principled approach yields sharp performance guarantees and enables a rigorous, apples-to-apples comparison between $\mathrm{EF}$, $\mathrm{EF^{21}}$, and compressed gradient descent. Our analysis is carried out in the simplified single-agent setting, which allows for clean theoretical insights and fair comparison of the underlying mechanisms.
Remark: proof certificates To consolidate and support our theoretical results, we complement each theoretical statement with analytical or numerical validation. Specifically, we provide certificates of correctness generated either with a Computer Algebra System (CAS), using a WolframScript, for symbolic verification, or using Performance Estimation Problems (PEP) for numerical validation. CAS enable verification of algebraic identities, while PEP annotations indicate numerical validation of complete statements. These certificates are highlighted in the paper using [Uncaptioned image] and [Uncaptioned image] markers, which are direct links to the corresponding Jupyter notebook or WolframScript in our public GitHub repository.  [^1] While these certificates do not replace the mathematical proofs presented in the paper, they serve as an additional layer of transparency and error checking, analogous to unit tests in software development. This practice provides a reproducible, independently verifiable basis for our theoretical claims, thereby reducing the risk of oversights in complex derivations.
@@ 1 Introduction
The compression operator $\mathcal{C}$ is a stochastic operator such that, for some $\epsilon\in[0,1)$,
$\text{for all}\quad x\in\mathbb{R}^{d},\quad\mathbb{E}\left[\|x-\mathcal{C}(x)\|^{2}\right]\leqslant\epsilon\|x\|^{2}.$ (2)
@@
More recently, a variant of the classic error feedback mechanism, known as $\mathrm{EF^{21}}$ was introduced by Richtárik et al. [14], and is presented in Algorithm 3. Unlike classic error feedback, $\mathrm{EF^{21}}$ focuses on communicating a gradient estimate that is more robust to the variance observed in gradients received from different agents around the minimum of finite sum objectives. This method has since been extended in several directions [e.g. 23, 36, 37, 38].
@@
To make this comparison sharp and transparent, we adopt a deliberately simple yet representative setup: we consider smooth and strongly convex functions in the single-agent setting ($n=1$). While simple, this regime is widely recognized as a crucial stepping stone—not only for building intuition, but also as its own theoretical contribution [e.g., 32, 10]. The single-agent setting is also of independent interest: in the field of sparsity-aware neural-network training, sparse-update methods have been shown to correspond to error feedback [43]. In this context, tightness means that we identify the best possible Lyapunov function within a given class and compute the exact worst-case convergence rate over the class of problems considered.
@@
This work makes the following contributions:
@@
    From a methodological perspective, this paper shows how to apply the performance estimation framework to algorithms from the federated learning literature that incorporate compression schemes. By leveraging this methodology—both analytically and numerically—this work paves the way for a more precise and reliable understanding of federated and distributed learning methods.
@@
    This work provides a tight analysis of $\mathrm{EF}$ and $\mathrm{EF^{21}}$, and compare them with compressed gradient descent in the single-agent setting, on $L$-smooth, $\mu$-strongly convex functions. In particular, we give an analytical formula of the best possible contraction rate, by analyzing an optimal Lyapunov function within a class of candidate Lyapunov functions defined in Definition 1. Furthermore, we provide the optimal tuning for the step size in both those algorithms.
@@
    We demonstrate that those rates are achieved, proving that the analysis is tight.
@@ 2.1 Theoretical results on $\mathrm{CGD}$, $\mathrm{EF}$, $\mathrm{EF^{21}}$
In the single agent case, we leverage the equivalence between compressed gradient descent ($\mathrm{CGD}$), under Assumption˜1 and the inexact gradient method with relatively bounded gradients. $\mathrm{CGD}$ corresponds to the particular case of Algorithm˜1 with $n=1$, and relatively bounded gradients means that for any $x$, the oracle queried at point $x$ outputs a value $g$ such that $\|g-\nabla f(x)\|^{2}\leq\epsilon\|\nabla f(x)\|^{2}$. Various notions of gradient approximation have been studied [49, 50, 51], and a tight analysis for relatively bounded gradients was given in [52]. Specifically, authors have shown that the inexact gradient method then enjoys tight convergence guarantees for any step size $\eta>0$, with respect to the functional residual, Euclidean norm distance to the solution and gradient norm. However, $\mathrm{CGD}$ is known to diverge when applied using stochastic gradient oracles, and to non-smooth functions [10]. Interestingly, it is also known to diverge in the multi-worker setting [24]. Studying $\mathrm{CGD}$ is important in its own right, because when the compression operator is chosen as the sign function, and the algorithm is applied in the stochastic setting (i.e., signSGD), there is a connection to Adam both in the convex [53], and non-convex setting [54].
Convergence rates for $\mathrm{EF}$ have been established with strongly convex [32], quasi-convex and non-convex [10, 35] functions, and using stochastic gradients. Richtárik et al. [14] study $\mathrm{EF^{21}}$ in the multi-worker setting, with (potentially) randomized compression operators. They establish a $\mathcal{O}(k^{-1})$ convergence rate on Lipschitz smooth functions, and a linear rate under the additional assumption that the functions satisfy the Łojasiewicz inequality. These results are obtained using a Lyapunov function; however, without tightness guarantees—neither for the choice of Lyapunov function, nor for the convergence rate itself. Extensions of $\mathrm{EF^{21}}$ have been proposed, including adaptations to stochastic gradients [23], and the introduction of a momentum term to improve sample complexity in the stochastic setting [36].
@@ 2.2 Performance estimation
Another relevant line of work we leverage to discover the analytical form of the Lyapunov functions lies in the field of symbolic regression, which aims to solve supervised learning tasks over the space of simple analytic expressions. Recent advances use genetic programming to search this space, and software packages have been developed in both Python and Julia [61].
@@ 2.3 Definitions & Notation
We have already introduced the notion of a contractive compression operator in Assumption˜1. To position our contribution within the broader literature we now specify that our analysis is restricted to the setting of smooth, strongly convex functions:
@@
The function $f$ is $L$-smooth, i.e., for all $x,y\in\mathbb{R}^{d}$, we have
$f(y)\leqslant f(x)+\langle\nabla f(x),y-x\rangle+\frac{L}{2}\|y-x\|^{2}.$
@@
The function $f$ is $\mu$-strongly convex, i.e., for all $x,y\in\mathbb{R}^{d}$, we have
$f(y)\geqslant f(x)+\langle\nabla f(x),y-x\rangle+\frac{\mu}{2}\|y-x\|^{2}.$
@@
We formally denote $\mathcal{M}:\mathbb{R}^{\ell\times d}\times\mathbb{R}^{d}\times\mathcal{F}\to\mathbb{R}^{\ell\times d}\times\mathbb{R}^{d}$ a first-order method acting on a set of functions $\mathcal{F}$, for an integer $\ell\in\mathbb{N}$. Such a method, given a function $f\in\mathcal{F}$, is applied to an initial state $\xi_{0}\in\mathbb{R}^{\ell\times d}$ and iterate $x_{0}\in\mathbb{R}^{d}$, and generates a sequence $\{\xi_{k}\}_{k\geqslant 0}$ of states, and a sequence $\{x_{k}\}_{k\geqslant 0}$ of iterations. The states represent information summarizing the current point in the optimization trajectory that the algorithms may depend on beyond the current iterate—for example, error-related quantities in error feedback algorithms. The integer $\ell$ is thus typically small, from 0 to 3 in general. The specific states used in this paper are all specified in Subsection˜3.1.
Definition 1 (Candidate Lyapunov function).
@@
    (Zero at fixed-point) $\mathcal{V}(\xi,x;f)=0$ if and only if $x=x_{\star}$ and $\xi=\xi_{\star}$ for a unique $\xi_{\star}\in\mathbb{R}^{\ell\times d}$.
@@
for some constant $\rho<1$ and for all $k\geqslant 0$, uniformly over $\mathcal{F}$. Finding the optimal Lyapunov function within a parameterized class, for a method $\mathcal{M}$, then amounts to solving the following problem:
@@
Note that we set $k=0$ in order to have a guarantee that is valid for all $x\in\mathbb{R}^{d}$.
@@ 3.1 Numerical performance of all methods
where all variables are defined as in Algorithm 1, Algorithm 2 and Algorithm 3, respectively. Note that all numerical results of this article are using deterministic compression operators.
@@
We are now ready to present the numerical results on the performance of each method. Figure 1 shows contour plots of the contraction factor for each method. That is, for a fine grid of both the step size $\eta$ in Algorithms˜1, 2 and 3, and the parameter $\epsilon$ in Assumption 1, we numerically compute the value of the best possible worst-case contraction rate in terms of our class of candidate Lyapunov functions, over $\mathcal{F}_{\mu,L}$, as given by (6). A darker blue point indicates a stronger contraction $\rho^{\star}(\mathcal{M})$ (i.e., a better rate). A red point indicates that the method is non-convergent for that choice of $(\epsilon,\eta)$. We observe that, although the results are purely numerical at this stage, they indicate that $\mathrm{EF}$ and $\mathrm{EF^{21}}$ exhibit identical performance in our setting. This numerical equivalence is supported by Table 2: the maximum absolute difference between contraction factors for $\mathrm{EF}$ and $\mathrm{EF^{21}}$ is on the order of $10^{-5}$ to $10^{-7}$. This first fact is a very surprising observation. Indeed while $\mathrm{EF}$ and $\mathrm{EF^{21}}$ are known to be identical in the very specific case of using a deterministic positively homogeneous and additive compression operator [14, see Section 4.2], $\mathrm{EF}$ and $\mathrm{EF^{21}}$ remain grounded in fundamentally different motivations at first sight: on the one hand, $\mathrm{EF}$ accounts for the errors introduced by the compression step, while on the other hand, $\mathrm{EF^{21}}$ subtracts a control variate from the gradient prior to compression. Proving that the best possible convergence rate they can obtain, for any tuning $\epsilon,\eta$, is similar (but achieved for a different Lyapunov function) was, to the best of our knowledge, never established in the literature. It thus constitutes a significant step towards better understanding their connections.
$\kappa=2$ $\kappa=4$ $\kappa=10$
Absolute error 4.34e-07 6.02e-07 1.21e-06
@@
$\kappa=2$ $\kappa=4$ $\kappa=10$
@@
Table 2: Maximum absolute difference of contraction factor for $\mathrm{CGD}$ when allowing any combination of terms in our Lyapunov function, compared to the contraction achieved by the functional residual. Same grid as Table˜2.
@@
While given for a single $(\mu,L)$ in Figure 1 and Table˜2, similar results hold for all values of $(\mu,L)$ that were tried numerically, and several examples are given in Appendix D, along with details of the numerical experiments, including the computation of regions of non-convergence.
Furthermore, we tune each algorithm by picking the optimal step size for each method. We compute the rate $\inf_{\eta}\rho^{\star}(\mathcal{M}_{\eta})$, for $\mathcal{M}\in\{\mathrm{CGD},\mathrm{EF},\mathrm{EF^{21}}\}$, where $\mathcal{M}_{\eta}$ corresponds to the method with step size $\eta$. Results are shown in Appendix˜B for three values of $\kappa$, namely 2, 4, and 10. In each setup, for every level of compression, $\mathrm{CGD}$ achieves a rate which is strictly better than $\mathrm{EF}$ and $\mathrm{EF^{21}}$. In Subsection˜B.1, we also provide a symbolic certifiate of this fact. These results challenge the prevailing intuition that error feedback ensures convergence comparable to that of uncompressed methods, and even demonstrate that in the single agent and deterministic gradient regime, error feedback is actually always detrimental to convergence.
Finally, we note that the functional residual constitutes an optimal Lyapunov function for $\mathrm{CGD}$, as shown in Table 2. This result is not particularly surprising, given that a tight analysis of the functional residual with the optimal step size was previously established for inexact gradient descent by De Klerk et al. [52]. That work also demonstrated the tightness of this rate [^2] Following the same line of reasoning as in our remark on the tightness of our Lyapunov functions in Section 5, we can then show that the functional residual is an optimal Lyapunov function for $\mathrm{CGD}$.. For $\mathrm{EF^{21}}$, Richtárik et al. [14] proposed another Lyapunov function than the one used here. In Subsection˜D.5, Figure˜12 we compare the complexities of an optimally tuned version of their Lyapunov function with the class-optimal Lyapunov function numerically. Further comparison of rates can be found in Appendix˜B. In Subsection˜B.2, we prove that our rate is strictly faster than the the aforementioned rate for $\mathrm{EF^{21}}$.
@@
3.2 Exact convergence rate and optimal tuning for $\mathrm{EF}$
We begin by stating the main result of this section, which is a tight rate of convergence for $\mathrm{EF}$.